Shift report 20000820 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sun Aug 20 2000 - 16:18:23 EDT

  • Next message: http account: "Shift report 20000820 16:00-24:00"

    ------------------------------------------------------------
    Date: Sun Aug 20 16:18:22 2000, Shift: 08:00-16:00
    Supervisor: ykl, with: b2
    Log book#4 pages 130 - 132
    
    Shift Summary:
    ------------------------------------------------------------
    1.Good beam most of the time. Beam was dumped once at 10:13
    and cogged and stored within 30 minutes.
    2. Run time was divided into 7 of one-hour runs from
    Run 2358 to run 2365 to accumulate approximately 350 K
    collisions. D5 at 2Kgause in Mode B.
    3. No additional detector problems beyond the existing
    problems with T1 and DC which were not on, and HV problems
    with BBL and H2. No attempt was made to fix them.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sun Aug 20 2000 - 16:21:44 EDT